40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 20 16:14:17 2017, gautam, Update, Equipment loan, Impedance test kit borrowed from Downs 
    Reply  Thu Sep 21 01:32:00 2017, gautam, Update, ALS, AUX X Innolight AM measurement running Innolight_AM.pdfInnolight_AM.tar.gzIMG_7599.JPG20170921_203741_TFAG4395A_21-09-2017_115547_FourSquare.pdf
       Reply  Thu Sep 21 01:55:16 2017, rana, Update, ALS, X End table of Shame 20170921_003143.jpg20170921_002430.jpg20170921_002243.jpg20170921_001906.jpg
          Reply  Thu Sep 21 15:23:04 2017, gautam, Omnistructure, ALS, Long cable from LSC->IOO 
          Reply  Fri Sep 29 11:16:52 2017, Steve, Update, ALS, Y End table corrected ETMYf1.jpg
          Reply  Fri Oct 6 17:08:09 2017, Steve, Update, ALS, X End table beam traps corrected c1.jpgc2.jpgc3.jpgc4.jpg
       Reply  Sun Sep 24 20:47:15 2017, rana, Update, Computer Scripts / Programs, RF TF Uncertainties TFAG4395A_21-09-2017_115547_FourSquare.pdfTFAG4395A_21-09-2017_115547.tgz
          Reply  Wed Sep 27 00:20:19 2017, gautam, Update, ALS, More AM sweeps TFAG4395A_26-09-2017_202344_FourSquare.pdflowTemp.tgzTFAG4395A_26-09-2017_231630_FourSquare.pdfhighTemp.tgz
             Reply  Wed Sep 27 23:44:45 2017, gautam, Update, ALS, Proposed PM measurement setup IMG_7609.JPG
Message ID: 13346     Entry time: Fri Sep 29 11:16:52 2017     In reply to: 13326
Author: Steve 
Type: Update 
Category: ALS 
Subject: Y End table corrected 

The first Faraday isolater rejected beam path from the NPRO is fixed.

 

Attachment 1: ETMYf1.jpg  221 kB  | Hide | Hide all
ETMYf1.jpg
ELOG V3.1.3-